OpenStudy (anonymous):

1. Select the proper equation in point-slope form for the given point and slope. Point: (1, 9); Slope: 5 (Points : 1) y – 1 = 5(x + 9) y – 9 = 5(x + 1) y – 9 = 5(x – 1) y + 9 = 5(x – 1) Question 2.2. Select the proper equation in point-slope form for the given point and slope. Point: (–8, 3); Slope: 6 (Points : 1) y + 3 = 6x – 48 y + 3 = 6(x + 8) y – 3 = 6(x – 8) y – 3 = 6(x + 8)

OpenStudy (anonymous):

point slope form is y-y1=m(x-x1) the point becomes (x1,y1) and m=slope

OpenStudy (anonymous):

so for the first one y-9=5(x-9) and the 2nd y-3=6(x-(-8) y-3=6(x+8)